OpenStudy (anonymous):

how do i graph 2x+y=7??

OpenStudy (torilovesvampires):

Step 1: Put the equation in Slope Intercept Form. Step 2: Graph the y-intercept point (the number in the b position) on the y-axis. ... Step 3: From the point plotted on the y-axis, use the slope to find your second point. ... Step 4: Draw your line using the two points you plotted (y-intercept (b) first, slope (m) second.

OpenStudy (torilovesvampires):

0 = -2X - 7 2X = -7 X = -7/2 Now plot (-7/2, 0) Then set X to 0 and solve for Y Y = -2(0) - 7 Y = -7 Now plot (0,-7) Run a line between those two points and you have your graph. I chose 0, but you can do this with any number. Try 1 1 = -2x -7 8 = -2x 8/-2 = x

OpenStudy (anonymous):

im still confused. :(

OpenStudy (anonymous):

wouldnt it be y=-2x+7? But how would i graph that?

OpenStudy (torilovesvampires):

point (0,7)

OpenStudy (anonymous):

So from (0,7) i just go down -2 and to the right 1?

OpenStudy (torilovesvampires):

yes

OpenStudy (anonymous):

and i continue doing that then thats my line?

OpenStudy (torilovesvampires):

Yes

OpenStudy (anonymous):

Thank you

OpenStudy (torilovesvampires):

Your Welcome